OpenStudy (anonymous):

If P (x+1)= 3x^2-2x-5 Find P(4).

OpenStudy (anonymous):

my answer is 35, i'm not sure

OpenStudy (anonymous):

Not quite. How did you get that?

OpenStudy (anonymous):

P(4)=3(4)^2-2(4)-5 3(16)-8-5 48-8-5 =35

OpenStudy (anonymous):

I dont know if I am right trying to ask for help here.

OpenStudy (anonymous):

There are actually several ways to do this... Easiest Way: Set x + 1 equal to 4 and solve for x. Then plug that in for x. x + 1 = 4 x = 3 Plug in 3 now.

OpenStudy (anonymous):

thanks for the help @Calcmathlete

OpenStudy (anonymous):

Do you understand? Right now, it's P(x + 1). It's asking for P(4). What value of x would make x + 1 = 4? x = 3.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

trying to solve now

OpenStudy (anonymous):

P(4)=16? @Calcmathlete

OpenStudy (anonymous):

I do believe so :)

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

np :)